OpenStudy (anonymous):

The result of a survey about what 432 students would like as a career are shown below. 9% lawyer 13% nurse 13% doctor 23% teacher 12% athlete 30% undecided 1. How many of the students want to be teachers? 2. How many students want to work as doctors or nurses? 3. How many students did not want to be lawyers?

OpenStudy (anonymous):

1. 0.23 * 432 = 99.36 2. 0.26 * 432 = 112.32 3. 432 (1 - 0.09) = 393.12

OpenStudy (anonymous):

how did you get that answer in 2

OpenStudy (anonymous):

13% doctors + 13% nurses = 26% ( doctors and nurses )

OpenStudy (anonymous):

but you don't have to add all together because it says doctor OR nurses

OpenStudy (anonymous):

how did you get number 3 i dont understand

OpenStudy (anonymous):

Notice that the preference to be a doctor or a nurse is %13 each. Together they add up to be 26%.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

how did you get number 3 i dont understand

OpenStudy (anonymous):

If 9% do not want to be lawyers, then 91% want to be something else other than a lawyer.